Solve by substitution.
7x-6y=-80
5x+67=y
What is the solution?
Type an ordered pair.

Answer by angelagallegos(8)

7x-6y=-80
5x+67=y
first you are going to put the second equation into y= format: y=5x+67
now plug in 5x+67 for y in the first equation
7x-6(5x+67)= -80 distribute the 6
7x-30x-402= -80 combine like terms
7x-30x= -23x
-23x-402= -80 get x by itself, divide both sides by -23
x=-14
Plug in -14 in for x in the first equation
y= 5(-14)+67
-70+67
y = -3
to check plug both x and y in first equation
-3=5(-14)+67
-3=-70+67
-3= -3 true
your point is (-14, -3)
hope this helps